OpenStudy (anonymous):

Which are complementary angles? A)35° and 65° B)40° and 50° C)30° and 65° D)15° and 165°

OpenStudy (amistre64):

complementary angles add up to equal 90

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

in a right triangle that is

OpenStudy (anonymous):

is it all was going to be in a right triangle?

OpenStudy (amistre64):

in every triangle there is only 180 degrees that you can use; in a right triangle 90 of them are already used up. which just leaves the other 90 to be divided between the other 2 angles...does that make sense?

OpenStudy (anonymous):

kinda

OpenStudy (amistre64):

3 angles in a triangle; lets name them A,B,C in "any" triangle the angles add up to 180 A + B + C = 180 with a right triangle one angle is already 90 degrees, lets say A = 90, we get 90 + B + C = 180 subtract each side by 90 to get B + C = 180-90 B + C = 90

OpenStudy (anonymous):

i get it now thanks

OpenStudy (amistre64):

:)